Is the ordered pair (0,5) a solution of the system of linear equations:
x + y = 5
-5x + 2y = 10

Found 2 solutions by Theo, actuary:
Answer by Theo(13342)

should be:

replace x with 0 and y with 5 in each equation to get:

      0 +  5 =  5
      0 + 10 = 10

Since both equations are true, then (0,5) is a solution to the system of equations.

Answer by actuary(112)

This question is asking you to substitute 0 for X and 5 for y in both equations. If the right hand value is equal to the left hand value after making the substitution, the answer is "yes". Otherwise, the answer is "no".
Substitute 0 for x and 5 for y in the first equation and then the second equation. Since 0 + 5 = 5 and -5*0 + 2*5 = 10 the answer is "yes".